# Pub-Guard-LLM: Detecting Fraudulent Biomedical Articles with Reliable Explanations

### v4

@@ 4.3 Fine-Tuning
The model is fine-tuned on $y$ to understand how to identify fraudulent articles via a binary classification task with:
@@
where $\lambda$ is a hyper-parameter that controls the balance between the two objectives. This learning process helps Pub-Guard-LLM acquire knowledge beyond simple binary labels, which results in strong performance with well-reasoned explanations.
Figure 2: Debate application mode for Pub-Guard-LLM
@@ 4.4 Application Modes
After fine-tuning, the model $f_{\theta}$ can be leveraged to identify fraudulent articles while providing explanations. To showcase its usability, we introduce three distinct application modes (see Appendix B.3 for input-output examples for each mode).
@@
$f_{\theta}(x,k)$ does not store all articles within its parameters, and, in some cases, an input article’s main argument may lack support from existing publications. This suggests that the article’s findings could be controversial and potentially fraudulent. To address this, we retrieve relevant top-$l$ articles $\mathcal{A}$ (all legitimate) to the input article $x$ from PubMed and use them to validate the claims made in $x$ by passing them as input into the model: $f_{\theta}(x,k,\mathcal{A})$ (see details in Appendix B.2).
@@
Recent research has shown that Multi-Agent Debate can enhance both factual accuracy and explainability in LLMs Liang et al. (2023); Du et al. (2023); Freedman et al. (2024). The core idea involves multiple agents presenting their arguments, while a judge oversees to reach a final decision.
To generate the debate process in our setting (Figure 2), we train two specialized Pub-Guard-LLMs: (1) a support reviewer $f_{\theta}^{s}$, trained on legitimate articles, generates supporting arguments $arg^{+}$; (2) an attack reviewer $f_{\theta}^{a}$, trained on retracted articles, generates attacking arguments $arg^{-}$. Thus, given input article $(x,k)$, the support reviewer $f_{\theta}^{s}$ consistently provides arguments in favor of the article while the attack reviewer $f_{\theta}^{a}$ presents counterarguments highlighting potential issues. Following this debate, we introduce a meta-reviewer $f_{\theta}^{meta}$ that is responsible for making the final decision based on the arguments presented by both reviewers. To train the meta-reviewer, we use a teacher model to generate debate-based explanations $\hat{e}$, which are then used to fine-tune $f_{\theta}^{meta}$ by reusing the multi-task learning in Equation 3. Overall, the debate-based detection framework is formulated as: $f_{\theta}^{meta}(x,k,arg^{+},arg^{-})$, where the meta-reviewer makes a decision by considering both the supporting and attacking arguments.
@@
Table 2: Performance in detecting fraudulent articles. We report average F1 scores over three seeds.
Test Breast Lung Ovarian Colorectal Avg
Vanilla 84.8 88.5 89.2 85.9 69.7 83.6
RAG 66.9 82.7 87.9 85.6 81.8 81.0
Debate 90.7 92.2 94.6 89.9 75.9 88.7
Table 3: Recall (average over 3 seeds) across modes
@@ 5.1 Baselines
(1) Fine-tuning Binary Classifiers. This method appends a classification head—a linear layer—on top of a language model to handle fraud detection. However, this method can only output binary labels (no explanations). The title, abstract, and metadata are concatenated using a special token, [SEP], and this combined textual input is fed into the classifier. Specifically, we use language models: SciBERT Beltagy et al. (2019), BioLinkBERT Yasunaga et al. (2022), Gatortron Yang et al. (2022), Llama-3.1 Dubey et al. (2024) and Bio-Llama ContactDoctor (2024). We also find that introducing additional models to represent metadata separately can be beneficial. Thus, we apply a lightweight short-text model, PEARL Chen et al. (2024), to represent metadata and obtain embeddings for authors, affiliations, and journal. We continue to use the original language model to represent the title and abstract. Subsequently, these embeddings are concatenated and fed into the classification head. We denote the resulting variant of language model X as “X + PEARL”.
(2) Few-shot Reasoning. LLMs are generally known as few-shot reasoners, meaning they can perform powerful reasoning with few learning examples Brown et al. (2020); Kojima et al. (2022). Compared to binary classifiers, LLMs can offer answers with explanations. We provide LLMs with a certain number of learning examples and prompt them to obtain binary answers with explanations. Here, we use the following instruction-aligned LLMs: Llama-3.1-Instruct Dubey et al. (2024), OpenScholar Asai et al. (2024), Bio-Llama ContactDoctor (2024) and PMC-Llama Wu et al. (2024).
@@ 5.2 Implementation Details
All approaches are implemented with PyTorch Paszke et al. (2019) and HuggingFace Wolf et al. (2020). We use Amazon EC2 g5.12xlarge instances with 4$\times$24 GiB for all experiments. The Teacher model used for distilling explanations is Mistral-Large MistralAI (2024) on Amazon Bedrock. Our model fine-tunes Llama-3.1-8B Dubey et al. (2024) using LoRA Hu et al. (2021) adapters (r=128, lora_alpha=128, lora_dropout=0.1). We train on the train set in the General partition for one epoch and AdamW 8-bit optimizer with a learning rate of 1e-4, batch size 4, gradient accumulation of 4 and the $\lambda$ is 1.
All baseline binary classifiers are fine-tuned with three different seeds: we report their performance based on the best validation set score. For few-shot reasoning, we set the sample size to zero (zero-shot reasoning) since adding contextual examples provides only minimal benefits (1–2 percentage points).
6 Results
6.1 Overall Performance
Table 2 gives the results of the comparison of Pub-Guard-LLM with the baselines.
First, we note that Pub-Guard-LLM, across all three modes, consistently outperforms the baselines. The performances across the three modes do not exhibit significant differences, but the RAG mode achieves the best score on average, which shows that introducing relevant articles is beneficial to verify the claim quality in abstracts. On the other hand, a key advantage of the debate mode is its notably higher recall, as shown in Table 3. Achieving high recall is crucial in the fraud detection task, as its primary goal is to flag suspected fraud and provide warnings to end-users.
Second, BERT-based models prove to be cost-effective while achieving competitive performance comparable to larger counterparts in domain-specific tasks, which aligns with prior findings Lehman et al. (2023); Chen and Varoquaux (2024). For instance, Gatortron + PEARL ranks second overall. Additionally, leveraging top-layer representations of decoder-only models to train classifiers appears suboptimal, as it does not perform better than BERT-based models. Existing LLMs struggle in the few-shot setting, suggesting that domain adaptation is necessary for this task. Note that we do not report the performance of GPT-4o and Mistral-Large because these large instruction-aligned models frequently classify all articles as legitimate to avoid controversial conclusions.
In summary, all three modes of Pub-Guard-LLM effectively detect fraudulent articles and each offers distinct advantages. The Vanilla Reasoning mode is efficient and user-friendly, the RAG mode achieves the best overall performance, and the Debate mode excels in recall. This variability makes Pub-Guard-LLM highly versatile, accommodating a wide range of possible user needs in fraudulent article detection. Additionally, we conduct an ablation study to validate the effect of each component in our system, such as base LLM, Teacher Model, External Knowledge. The results are in Table A1 in the Appendix.
Figure 3: Explanation reliability using LLM-as-a-Judge (average over three runs)
@@ 6.1 Overall Performance
Figure 5: Comparison of explanation reliability and label-explanation agreement in the user study (see details in Appendix C.2).
@@ 6.2 Explanation Reliability
Compared to BERT-based models, a distinct benefit of Pub-Guard-LLM is the explanations. To evaluate the reliability of LLM-generated explanations, we consider two dimensions introduced, for different settings, in prior work Kotonya and Toni (2020, 2024); Valentino and Freitas (2024). (1) Relevance. A relevant explanation should highlight the particular features and patterns within the article that the model identified as indicative of fraud or not. The model should only reference information that is present in the article, avoiding hallucinations. (2) Coherence. This refers to the logical consistency and clarity of the explanation provided by the model for its predictions. Coherence ensures that the explanation flows logically, thereby enhancing the user’s understanding of why an article is classified as fraudulent or not. To measure these properties, we adopt the LLM-as-a-Judge paradigm, which has been shown to align with both controlled and crowdsourced human preferences when using strong LLM judges like GPT-4  Zheng et al. (2023). We design prompts (see Appendices A4 and A5) for these two dimensions and ask GPT-4o to assess the provided explanations, which assigns a score between 1 and 10. The results, presented in Figure 3, show that our explanations outperform those of other LLMs, with the exception of OpenScholar. While OpenScholar demonstrates strong explanation quality, its fraud detection capability is significantly lower than other LLMs, with an average F1 score of just 16.9 (see Table 2). Also, note that the debate mode can reduce irrelevant explanations by achieving higher relevance scores than the another two modes. These findings validate that Pub-Guard-LLM not only achieves powerful performance but also provides reliable explanations.
6.3 Impact of Metadata Levels
We explore whether the external knowledge we use to assess levels of the metadata can provide useful heuristics for fraud detection. To test this, we compare Pub-Guard-LLM (Vanilla) to heuristic methods, e.g., using the credibility of authors as indicators (see Appendix C.1). The results are shown in Figure 4. First, we observe that relying solely on heuristic features is insufficient for fraud detection since Pub-Guard-LLM significantly outperforms heuristics on average. Second, prestigious researchers, institutions and journals should not be blindly trusted, and these high-profile frauds are harder to detect, as shown by the performance declines on high and very high groups in Figure 4a and Figure 4b. Additionally, null is an indicator of low-level credibility and Pub-Guard-LLM is robust against missing information. In summary, high-profile cases pose a significant challenge to fraud detection. In the future, we plan to investigate features beyond metadata such as images.
6.4 User Study
@@
We conducted a pilot user study to evaluate whether the outputs of Pub-Guard-LLM align with users’ needs. To achieve this, we designed questionnaires covering our three application modes and gathered feedback from three experienced doctors (see details in Appendix C.2). The questionnaire focused on two key aspects: (1) Reliability: is the explanation coherent and relevant? (2) Agreement: do users agree with the predicted label and the explanation? Figure 5 presents the user study results. The feedback suggests that doctors find our model’s explanations reliable, but there is a discrepancy between expert opinions and machine-generated predictions regarding the debate mode. However, due to the limited number of participants, it is difficult to determine which application mode performs better.
@@ 7 Conclusion
In this work, we contribute to fraudulent academic articles by introducing a publicly available benchmark and the first open-source LLM-based system tailored for this task. Our Pub-Guard-LLM not only surpasses existing baselines but also delivers reliable, well-grounded explanations. We aim to raise awareness of academic fraud within the research community and drive further advancements to tackle this growing issue.
We emphasize the need for broader efforts to combat academic fraud in the following key areas. First, high-profile publications should not be blindly trusted, as fraudulent publications may still exist within them. Detecting such cases requires more than analyzing metadata and abstracts, and future studies should also investigate full texts and images to identify frauds. Second, as we develop technological solutions to detect fraud, entities like paper mills continuously improve their methods to circumvent detection. This ongoing arms race requires continuous adaptation of detection techniques. Finally, addressing scientific fraud fundamentally depends on human and societal factors. Researchers should adhere to scientific integrity, and our academic evaluation system should prioritize genuine value creation over mere quantitative metrics.
@@
While our study makes contributions to fraudulent article detection, we acknowledge two primary limitations: (1) Catastrophic forgetting. Pub-Guard-LLM is a task-specific model that may not perform well in general-purpose reasoning due to catastrophic forgetting McCloskey and Cohen (1989)—a phenomenon where a neural model loses previously acquired knowledge upon learning new information. (2) Limited participants in user study. Our user study involved a small number of participants, which limits the ability to systematically compare the three application modes. A more comprehensive study with a larger participant pool is necessary for thorough evaluation.
@@ Appendix A Details of Benchmark and External Knowledge
[…]
@@
  • World-Class Institution (very high): average citations $>45$
@@
[…]
@@ B.1 Prompts
Here, we list the prompts used in this work: (1) Explanation Distillation, Table A2; (2) Prompt of Detecting Fraudulent Articles, Table A3; (3) Prompt of Relevance Evaluation, Table A4; (4) Prompt of Relevance Coherence, Table A5.
@@ B.2 RAG Mode
We retrieve relevant top-$l$ articles $\mathcal{A}$ (all legitimate) to the input article $x$ from PubMed and use them to validate the claims made in $x$. In our experiment, the value $l$ is 5 and the PubMed database is MedRAG PubMed [^3] https://huggingface.co/NeuML/pubmedbert-base-embeddings with 3.5M articles. The retriever here is PubMedBERT and we use cosine similarity to compute the semantic relatedness.
@@ C.1 Heuristic Features
Since we incorporate external knowledge to represent metadata levels, these fine-grained features can serve as indicators for fraud detection. To leverage this, we design three heuristics based on different metadata attributes. The intuition is that articles from prestigious researchers or well-known institutions are generally perceived as more trustworthy. To classify articles as fraudulent or not, we establish a threshold-based approach using metadata credibility scores. For example, if an article’s highest author credibility falls within [medium, high, very high], we classify it as legitimate. Author credibility and affiliation reputation are categorized into six levels (including null). We define the three lower levels as indicators of fraudulence and the three higher levels as indicators of legitimacy. Journal reputation, which consists of five levels, is categorized similarly: the three lower levels indicate potential fraud, while the higher levels suggest legitimacy.
C.2 User Study
In the user study, we prepared three questionnaires to evaluate the Vanilla Reasoning, RAG and Debate modes independently. Using the Vanilla Reasoning mode, we generated explanations for 300 papers. We randomly selected five fraudulent papers and five legitimate papers from the result and again randomized the order of the paper to create the questionnaire. For each paper, we added 1) the information of the paper as it was in the model prompt; and 2) the explanations generated by the model. Three questions were asked for each paper: 1) whether the explanation was coherent and relevant; 2) to infer the predicted label from the given explanation; and 3) whether the user agreed with the actual prediction. The same process was followed for the other two questionnaires for the Debate and RAG models, except that, for the Debate model, the information of a supporting viewer and an attacking reviewer was added to the paper description. In total, 30 distinct papers were selected for evaluation by three clinicians.

### v5

@@ 4.3 Fine-Tuning
The model is fine-tuned on $y$ to understand how to identify retracted articles via a binary classification task with:
@@
where $\lambda$ is a hyper-parameter that controls the balance between the two objectives. This learning process helps Pub-Guard-LLM acquire knowledge beyond simple binary labels, which results in a strong performance with well-reasoned explanations.
@@ 4.4 Application Modes
After fine-tuning, the model $f_{\theta}$ can be leveraged to identify retracted articles while providing explanations. To showcase its usability, we introduce three distinct application modes (see Appendix B.3 for input-output examples for each mode).
@@
$f_{\theta}(x,k)$ does not store all articles within its parameters, and, in some cases, an input article’s main argument may lack support from existing publications. This suggests that the article’s findings could be controversial and potentially retracted. To address this, we retrieve relevant top-$l$ articles $\mathcal{A}$ (all legitimate) to the input article $x$ from PubMed and use them to validate the claims made in $x$ by passing them as input into the model: $f_{\theta}(x,k,\mathcal{A})$ (see details in Appendix B.2).
@@
Recent research has shown that Multi-Agent Debate can enhance both factual accuracy and explainability in LLMs (Liang et al., 2023; Du et al., 2023; Freedman et al., 2024). The core idea involves multiple agents presenting their arguments, while a judge oversees to reach a final decision.
Figure 2: Debate application mode for Pub-Guard-LLM
@@
Table 2: Performance in detecting retracted articles. We report average F1 scores over three seeds.
To generate the debate process in our setting (Figure 2), we fine-tune three specialized Pub-Guard-LLMs following the steps in Section 4.3: (1) a support reviewer $f_{\theta}^{s}$, trained on legitimate articles, generates supporting arguments $arg^{+}$; (2) an attack reviewer $f_{\theta}^{a}$, trained on retracted articles, generates attacking arguments $arg^{-}$. Thus, given input article $(x,k)$, the support reviewer $f_{\theta}^{s}$ consistently provides arguments in favor of the article while the attack reviewer $f_{\theta}^{a}$ presents counterarguments highlighting potential issues. Following this debate, we introduce (3) a meta-reviewer $f_{\theta}^{meta}$ that is responsible for making the final decision based on the arguments presented by both reviewers. To train the meta-reviewer, we use a teacher model to generate debate-based explanations $\hat{e}$, which are then used to fine-tune $f_{\theta}^{meta}$ by reusing the multi-task learning in Equation 3. Overall, the debate-based detection framework is formulated as: $f_{\theta}^{meta}(x,k,arg^{+},arg^{-})$, where the meta-reviewer makes a decision by considering both the supporting and attacking arguments.
@@ 5.1 Baselines
(1) Fine-tuning Binary Classifiers. This method appends a classification head—a linear layer—on top of a language model to handle retraction detection. However, this method can only output binary labels (no explanations). The title, abstract, and metadata are concatenated using a special token, [SEP], and this combined textual input is fed into the classifier. Specifically, we use language models: SciBERT (Beltagy et al., 2019), BioLinkBERT (Yasunaga et al., 2022), Gatortron (Yang et al., 2022), Llama-3.1 (Dubey et al., 2024) and Bio-Llama (ContactDoctor, 2024). We also find that introducing additional models to represent metadata separately can be beneficial. Thus, we apply a lightweight short-text model, PEARL (Chen et al., 2024), to represent metadata and obtain embeddings for authors, affiliations, and journal. We continue to use the original language model to represent the title and abstract. Subsequently, these embeddings are concatenated and fed into the classification head. We denote the resulting variant of language model X as “X + PEARL”.
(2) Zero-shot Reasoning. Compared to binary classifiers, LLMs can offer answers with explanations by using prompt-based methods Kojima et al. (2022). We provide LLMs with a dedicated prompt to obtain binary answers with explanations. Here, we use the following instruction-aligned LLMs: Llama-3.1-Instruct (Dubey et al., 2024), OpenScholar (Asai et al., 2024), Bio-Llama (ContactDoctor, 2024) and PMC-Llama (Wu et al., 2024).
@@ 5.2 Implementation Details
[…]
All baseline binary classifiers are fine-tuned with three different seeds: we report their performance based on the best validation set score.
@@
6 Results
6.1 Overall Performance
Table 2 gives the results of the comparison of Pub-Guard-LLM with the baselines. First, we note that Pub-Guard-LLM, across all three modes, consistently outperforms the baselines. The performances across the three modes do not exhibit significant differences, but the RAG mode achieves the best score on average, which shows that introducing relevant articles is beneficial to verify the claim quality in abstracts. On the other hand, a key advantage of the debate mode is its notably higher recall, as shown in Appendix Table A1. Achieving high recall is crucial in the retraction detection task, as its primary goal is to flag potential retraction and provide warnings to editors. Second, BERT-based models prove to be cost-effective while achieving competitive performance comparable to larger counterparts in domain-specific tasks, which aligns with prior findings (Lehman et al., 2023; Chen and Varoquaux, 2024). For instance, Gatortron + PEARL ranks second overall. Additionally, leveraging top-layer representations of decoder-only models to train classifiers appears suboptimal, as it does not perform better than BERT-based models. Existing LLMs struggle in the zero-shot setting, suggesting that existing LLMs do not have sufficient knowledge for this task. Note that we do not report the performance of GPT-4o and Mistral-Large because these large instruction-aligned models frequently classify all articles as legitimate to avoid controversial conclusions.
In summary, all three modes of Pub-Guard-LLM effectively detect retracted articles and each offers distinct advantages. The Vanilla Reasoning mode is efficient and user-friendly, the RAG mode achieves the best overall performance, and the Debate mode excels in recall. This variability makes Pub-Guard-LLM highly versatile, accommodating a wide range of possible user needs in retracted article detection. Additionally, we conduct an ablation study to validate the effect of each component in our system, such as base LLM, Teacher Model, External Knowledge. The results are in Table A2 in the Appendix.
6.2 Pub-Guard-LLM and Metadata Bias
Our proposed method relies on metadata like authors and affiliations. To test whether this metadata introduces biases, specifically against junior researchers or less-known universities, we compare Pub-Guard-LLM (Vanilla) to heuristic methods, which use only individual metadata information to classify retractions and are thus biased. Fo example, we use, as a heuristic method, the credibility level of authors to predict whether a given article should be retracted (see Appendix C.1).
The results are shown in Figure 3. The red horizontal line in each subfigure represents the F1 score achieved by the heuristic methods (i.e. metadata alone). Our findings indicate that Pub-Guard-LLM does not over-rely on a single metadata component. Rather, it integrates both abstracts and three metadata features to make final decisions. This finding proves that Pub-Guard-LLM can mitigate the metadata bias. Furthermore, prestigious researchers, institutions and journals should not be blindly trusted, and these high-profile retractions are harder to detect, as shown by the performance declines on high and very high groups in Figure 3a and Figure 3b. Note that the null is an indicator of low-level credibility and Pub-Guard-LLM is robust against missing information. In summary, high-profile cases pose a significant challenge to retraction detection.
(a) Explanation reliability using LLM-as-a-Judge (average over three runs)
(b) Comparison of explanation reliability and label-explanation agreement in the user study (see details in Appendix C.3)
Figure 4: Explanation evaluation and user study
@@ 6.3.1 Automatic Evaluation
Compared to BERT-based models, a distinct benefit of Pub-Guard-LLM is the explanations. To evaluate the reliability of LLM-generated explanations, we consider two dimensions introduced, for different settings, in prior work (Kotonya and Toni, 2020, 2024; Valentino and Freitas, 2024). (1) Relevance. A relevant explanation should highlight the particular features and patterns within the article that the model identified as indicative of retraction or not. The model should only reference information that is present in the article, avoiding hallucinations. (2) Coherence. This refers to the logical consistency and clarity of the explanation provided by the model for its predictions. Coherence ensures that the explanation flows logically, thereby enhancing the user’s understanding of why an article is classified as retracted or not. To measure these properties, we adopt the LLM-as-a-Judge paradigm, which has been shown to align with both controlled and crowdsourced human preferences when using strong LLM judges like GPT-4  (Zheng et al., 2023). We design prompts (see Appendices A5 and A6) for these two dimensions and ask GPT-4o to assess the provided explanations, which assigns a score between 1 and 10. The results, presented in Figure 4a, show that our explanations outperform those of other LLMs, with the exception of OpenScholar. While OpenScholar demonstrates strong explanation quality, its retraction detection capability is significantly lower than other LLMs, with an average F1 score of just 16.9 (see Table 2). Also, note that the debate mode can reduce irrelevant explanations by achieving higher relevance scores than the another two modes. These findings validate that Pub-Guard-LLM not only achieves powerful performance but also provides reliable explanations.
6.3.2 User Study
@@
We conducted a pilot user study to evaluate whether the outputs of Pub-Guard-LLM align with users’ needs. To achieve this, we designed questionnaires covering our three application modes and gathered feedback from three experienced doctors (see details in Appendix C.3). The questionnaire focused on two key aspects: (1) Reliability: is the explanation coherent and relevant? (2) Agreement: do users agree with the predicted label and the explanation?
Figure 4b presents the user study results. The feedback suggests that doctors find our model’s explanations reliable, but there is a discrepancy between expert opinions and machine-generated predictions regarding the debate mode. However, due to the limited number of participants, it is difficult to determine which application mode performs better.
@@ 7 Conclusion
In this work, we introduce a publicly available benchmark and the first LLM-based system tailored for retraction detection. Our Pub-Guard-LLM not only surpasses existing baselines but also delivers reliable, well-grounded explanations. We aim to provide open-source datasets and LLMs for this task, which helps stop retracted articles early and mitigates misinformation spread.
We emphasize the need for broader efforts to detect retractions in the following key areas. First, high-profile publications should not be blindly trusted, as retracted publications may still exist within them. Second, our method rely on metadata, which may introduce bias against junior researchers and new institutions. Detecting the above cases requires more than analyzing metadata and abstracts, and future studies should also investigate full texts, tables and images to identify retraction.
@@
While our study makes contributions to retracted article detection, we acknowledge several limitations: (1) Catastrophic forgetting. Pub-Guard-LLM is a task-specific model that may not perform well in general-purpose reasoning due to catastrophic forgetting (McCloskey and Cohen, 1989)—a phenomenon where a neural model loses previously acquired knowledge upon learning new information. (2) Limited participants in user study. Our user study involved a small number of participants, which limits the ability to systematically compare the three application modes. A more comprehensive study with a larger participant pool is necessary for thorough evaluation. (3) Metadata bias. Our method may create bias against researchers, journals and institutions, as suggested by our user study (Section C.3). (4) Absence of nuance in classification. The current dataset offers only binary labels, retracted or not. Future work should refine labels and break down retraction into different reasons, such as scientific misconduct, ethical violations, and honest error.
@@ Appendix A Details of Benchmark and External Knowledge
  • Emerging Researcher (very low): $0\leq h\text{-index}\leq 5$
  • Early Career Researcher (low): $6\leq h\text{-index}\leq 15$
  • Established Researcher (medium): $16\leq h\text{-index}\leq 30$
  • Influential Researcher (high): $31\leq h\text{-index}\leq 45$
  • Leading Expert (very high): $h\text{-index}>45$
@@
  • World-Class Institution (very high): average citations >45>45
@@
We define an article as high-profile if its metadata includes at least one leading expert (very high) author, or one world-class (very high) affiliation, or is published in a top-level (very high) journal. This definition aligns with our intuition, as articles produced by renowned researchers, prestigious institutions, or highly reputable venues are generally perceived as more trustworthy. We compute the high profile rate of retraction in each subset by: $|\textit{High Profile Articles}|/|\textit{Retracted Articles}|$. These high-profile cases are hard to detect, which makes the task particularly challenging.
@@ B.1 Prompts
Here, we list the prompts used in this work: (1) Explanation Distillation, Table A3; (2) Prompt of Detecting Retracted Articles, Table A4; (3) Prompt of Relevance Evaluation, Table A5; (4) Prompt of Relevance Coherence, Table A6.
@@ B.2 RAG Mode
We retrieve relevant top-$l$ articles $\mathcal{A}$ (all legitimate) to the input article $x$ from PubMed and use them to validate the claims made in $x$. In our experiment, the value $l$ is 5 and the PubMed database is MedRAG PubMed [^3] https://huggingface.co/datasets/MedRAG/pubmed with 3.5M articles. The embedding model here is PubMedBERT [^4] https://huggingface.co/NeuML/pubmedbert-base-embeddings and we use cosine similarity to compute the semantic relatedness.
@@ C.1 Heuristic Features
Since we incorporate external knowledge to represent metadata levels, these fine-grained features can serve as indicators for retraction detection. To leverage this, we design three heuristics based on different metadata attributes. The intuition is that articles from prestigious researchers or well-known institutions are generally perceived as more trustworthy. To classify articles as retracted or not, we establish a threshold-based approach using metadata credibility scores. For example, if an article’s highest author credibility falls within [medium, high, very high], we classify it as legitimate. Author credibility and affiliation reputation are categorized into six levels (including null). We define the three lower levels as indicators of retraction and the three higher levels as indicators of legitimacy. Journal reputation, which consists of five levels, is categorized similarly: the three lower levels indicate potential retraction, while the higher levels suggest legitimacy.
C.2 Evaluation of Distilled Explanation
In this evaluation, we provide our co-authors (NLP people and medics) with questionnaires. This questionnaire is designed to evaluate the quality of explanations generated by an LLM. For each case, the LLM receives information about a medical publication and generates an explanation according to the "Is Retracted" label. This evaluation focuses on this key question “ Is the explanation coherent and relevant? Coherent means the explanation is logical and consistent. Relevant means the explanation is closely connected and appropriate to the provided article information.” For each article, our co-authors answer one binary question and can add optional comments if they wish. Finally, we evaluate 45 randomly selected articles and 71.1% of these explanations are coherent and relevant.
C.3 User Study
In the user study, we prepared three questionnaires to evaluate the Vanilla Reasoning, RAG and Debate modes independently. Using the Vanilla Reasoning mode, we generated explanations for 300 papers. We randomly selected five retracted papers and five legitimate papers from the result and again randomized the order of the paper to create the questionnaire. For each paper, we added 1) the information of the paper as it was in the model prompt; and 2) the explanations generated by the model. Three questions were asked for each paper: 1) whether the explanation was coherent and relevant; 2) to infer the predicted label from the given explanation; and 3) whether the user agreed with the actual prediction. The same process was followed for the other two questionnaires for the Debate and RAG models, except that, for the Debate model, the information of a supporting viewer and an attacking reviewer was added to the paper description. In total, 30 distinct papers were selected for evaluation by three clinicians.